The length of a rectangle is 4 inches more than three times the width. The perimeter is 96 inches. Find the length and width.
Answer by mananth(16946)

let width be x
the length = 3x+4
..
2L +2W= perimeter
2x+2(3x+4)= 96
2x+6x+8=96
8x+8=96
-8
8x=96-8
8x=88
/8
x=11 in. the width
..
length = 3x+4
=3*11 +4
=33+4
=37 in the length
...
m.ananth@hotmail.ca